smatteson4768:

Kareem owes $5,127 on a credit card with a 15.9% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card?

darkknight:

\[A = P(1+r/n)^(nt)\] Where r = rate (15.9%) n= number of times compounded (12) A = Final investment ($5127) t = time (12 months)